What is the solution to 2*2^x+1 -5=15
Answer by lwsshak3(6494)

What is the solution to 2*2^(x+1) -5=15
..
2*2^(x+1) -5=15
2*2^(x+1) =20
2^(x+1)=20/2=10
(x+1) log2=log 10=1
xlog2+log2=1
xlog2=1-log2
x=(1-log2)/log2
x=2.3219 (ans)
..
Check
2*2^(x+1) -5
2*2^(2.3219+1)-5
2*2^(3.3219)-5=14.999..